# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

### v3

@@ Abstract
This paper presents a minimal operator-only term rewriting system (KO7) designed to investigate the foundations of self-verifying computation. The system consists of seven constructors and eight reduction rules, with no external axioms, no imported logic, and no pre-existing arithmetic.
Standard termination methods fail on this system. We document ten distinct proof strategies that succumb to the system’s structural complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, and unchecked recursion. These methods collapse at the same barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats every additive measure ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). This failure is not a defect of the proof search but a necessary feature of any system capable of sequential computation.
To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit. For this fragment, a novel triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and ordinal ranking) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
We state a fundamental conjecture: No relational operator-only TRS can have its full-system termination proved by internally definable methods. We define a “relational” TRS as one capable of internalized sequential computation. This property is shared by Turing machines (via transition functions), lambda calculus (via fixed-point combinators), and recursive functions. All such systems require a recursor to define the “next” step; this recursor inherently creates the duplication barrier.
This result reveals a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), this barrier appears at a more primitive layer: in operator systems possessing only order and recursion, before arithmetic emerges. The system’s inability to prove its own termination from within is a structural limitation of ordered computation itself, paralleling the halting problem but manifesting at the substrate of operator definitions.
The complete Lean 4 formalization ( 7,000 lines) is available at https://github.com/MosesRahnama/OperatorKO7.
@@ 1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture regarding the unprovability of termination for relational operator-only systems using internal methods.
Scope. The formal results apply to a guarded safe subrelation. This restriction is necessary because the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ when $\kappa^{M}(a)=0$. The safe fragment employs guarded/context joiners to ensure unique normal forms (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
@@
Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. This result supports a certified normalizer with proven totality and soundness. Furthermore, confluence for the safe fragment is established via Newman’s Lemma given verified local confluence. The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
@@
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
@@
  • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
  • Full Step caveat: A specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails is exhibited, justifying the restriction to the safe subrelation.
  • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
2 Background: TRSs, SN, reachability, and Newman
Standard abstract reduction and term rewriting notions apply [2, 8]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [7], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [1]. Modularity holds in certain linear, non-collapsing combinations [3], while even flat non-linear systems exhibit undecidable reachability and confluence [6]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [4] or (recursive) path orders [5].
@@ 3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing.
@@
  • Equality (meta-level consequence under confluence):
    • If $\mathrm{nf}(a)=\mathrm{nf}(b)$ then $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{void}$.
    • Otherwise $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, confluence ensures these outcomes are unique.
@@ 4 Strong normalization
The termination argument relies on a triple-lexicographic measure
$\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are formally handled (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). Duplicating branches use a compact MPO-style head precedence.
Step vs SafeStep (measure unification)
For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices. Steps in the full relation are covered by a disjunctive decrease certificate: each step corresponds to either a KO7 lex drop or an MPO triple drop.
Theorem 1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}(t)$.
@@
DM vs MPO on rules (explicit)
For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$. For eqW, we use a compact MPO-leaning measure where the head precedence orients $\mathrm{eqW}\,a\,b$ strictly above $\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
  • Base-order premise (merge-cancel): in $\mathrm{merge}\,t\,t\to t$, the RHS is strictly smaller than the LHS.
  • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
@@
Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
@@
so there is no strict drop when $\rho(S)\geq 1$. The robust fix uses DM/MPO: replace one element by a multiset of strictly smaller elements (DM), or use RPO/MPO with a precedence/status such that the LHS redex strictly dominates each RHS piece.
Corollary 1 (Strong normalization).
@@
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 1).
@@
The fix is DM/MPO: ensure each RHS piece is strictly smaller than the removed redex in a base order and lift via a multiset/path extension.
@@ 5 A certified normalizer
A normalization function is defined by well-founded recursion on $\mu^{3}$. Properties proven in the formalization include:
(Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
(Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
@@ 6 Local confluence and Newman (guarded safe relation)
Local confluence is discharged by joining the finite set of critical pairs. Combining Cor. 1 with Newman’s Lemma [7] yields:
Theorem 2 (Confluence and unique normal forms).
@@
Instantiation (SafeStep). Combining Cor. 1 (SN for SafeStep) with local-join lemmas yields confluence and unique normal forms for the safe fragment by Newman’s Lemma. In contrast, the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ under $\kappa^{M}(a)=0$; thus full confluence does not hold. The SafeStep relation restricts eqW to cases where arguments are guarded, preventing this divergence.
@@
  • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
@@
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which is shown to be not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
@@
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ not locally joinable at root
Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
@@
Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$. Here $\mathrm{deltaFlag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
@@
$\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
7 Impossibility results: The Ten Failed Methods
The formal development establishes the failure of ten distinct termination strategies, necessitating the use of the triple-lexicographic measure with Dershowitz-Manna orderings. These failures are not merely heuristic; they are formally witnessed by counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Operational_Incompleteness.lean).

### v4

# Strong Normalization for the Safe Fragment of a Minimal Rewrite Calculus: A Triple-Lexicographic Proof and a Conjecture on Full-Termination Limits for Pure Recursive Calculi (PRC)

@@ Abstract
This paper introduces the concept of a Pure Recursive Calculus (PRC), a minimal computational substrate defined by four strict structural conditions: no variable binding, no external memory, no imported axioms, and an unrestricted step-duplicating recursor. In our PRC framing, this recursor is the mechanism used to represent sequential iteration. We present KO7, a concrete PRC consisting of seven constructors and eight reduction rules, designed to investigate the foundations of self-verifying computation.
Standard termination methods fail on this calculus due to a core structural barrier: the duplication of the step argument in the recursor. The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats additive measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$). Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth. Furthermore, classical path orders (LPO/MPO) bypass this barrier by relying on the universal Subterm Property, which acts as an imported external axiom. Counting these external path-order methods yields fourteen total strategy families (Section 7). This pattern indicates a structural limitation for calculi of this form rather than a defect of proof search.
To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit. For this fragment, a triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and a computable natural rank) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
We state a central conjecture: For Pure Recursive Calculi (PRC), no internally definable method currently known to us proves full-calculus termination. Unlike Turing machines (which require an external tape), lambda calculus (which requires variable binding), and recursive functions (which import zero and successor as external primitives), a PRC has none of these external scaffolds. On this framing, the termination barrier is positioned at a more primitive layer than those classical settings.
These findings motivate a form of Pre-Arithmetic Incompleteness. Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), the barrier studied here appears at a more primitive layer: in operator calculi possessing only order and recursion, before arithmetic emerges. In this setting, the difficulty of proving full termination from internal methods parallels halting-style limitations at the substrate of operator definitions.
The complete Lean 4 formalization ($\sim$6,000 LOC) is available at https://github.com/MosesRahnama/OperatorKO7.
@@ 1 Introduction
This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules. There are no binders, types, external axioms, or semantic predicates; the rules are the semantics. The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the unprovability of termination for Pure Recursive Calculi (PRC) using internal methods.
Scope. The formal results apply to a guarded safe subrelation. This restriction is necessary because the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$. The safe fragment employs guarded/context joiners; the Newman local-join hypothesis is explicitly formulated and fully discharged in the artifact (§6).
Why KO7?
[…]
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi. If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization. With confluence, decision reduces to normal-form equality.
@@
Strong normalization is proven for a guarded fragment of KO7 using a triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and a computable natural rank. This result supports a certified normalizer with proven totality and soundness. Furthermore, a Newman engine for the safe fragment is formalized, and the global local-join hypothesis is fully discharged, yielding unconditional confluence and unique normal forms for the safe fragment.
Furthermore, we apply established proof theory to classify the boundaries of internal termination proving. Drawing on the work of Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi. These methods achieve termination by introducing the universal Subterm Property ($f(t)>t$) as a precondition, which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension). Because these orders import external structural axioms and well-ordering principles, they are outside the internal method class used in this paper. While not a proof of general non-termination, the work includes formally verified impossibility results for fourteen distinct strategy families under duplication. These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
@@
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and a computable natural rank.
@@
  • Newman (SafeStep): A Newman theorem is formalized for SafeStep; the global local-join hypothesis is discharged in the artifact, yielding unconditional confluence and unique normal forms for the safe fragment.
  • Full Step caveat: A concrete non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is exhibited for the full relation, justifying restriction to a guarded fragment.
  • Impossibility results: Twelve distinct impossibility barriers (including additive, polynomial, weight-based, and precedence-based measures) are formally machine-checked in Impossibility_Lemmas.lean and Conjecture_Boundary.lean, alongside two meta-theoretical barriers.
2 Background: Term Rewriting, SN, reachability, and Newman
Standard abstract reduction and term rewriting notions apply [2, 13]. A term is in normal form if no rule applies. A rewrite calculus is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [11], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating calculi has well-charted complexity: NP-complete for length-reducing calculi (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating calculi [1]. Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear calculi exhibit undecidable reachability and confluence [10]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [5] or (recursive) path orders [7].
@@ 3 The KO7 calculus
KO7 is a finite rewrite calculus over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified SafeStep fragment adds guards: R_eq_diff requires $a\neq b$, R_merge_void_left/right require $\delta$-flag tie conditions, R_merge_cancel requires both $\delta$-flag and $\kappa^{M}$ guards, R_rec_zero requires a $\delta$-flag guard on $b$, and R_eq_refl requires $\kappa^{M}(a)=0$. R_merge_cancel is collapsing. (* Note: R_rec_succ duplicates the step argument $s$ (one occurrence on the LHS, two on the RHS). This duplication is the termination stressor: the paper proves strong normalization for the guarded SafeStep fragment, not for the unrestricted full relation.)
@@
  • Strict Syntactic Equality:
    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (provided $\kappa^{M}(a)=0$).
    • If $a\neq b$, then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, arguments inside $\mathrm{eqW}$ are not subject to context closure, enforcing strict syntactic comparison rather than normal-form evaluation.
@@ 4 Strong normalization
The canonical termination argument relies on a computable triple-lexicographic measure
$\mu^{3}_{c}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\tau(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna); and (iii) a computable natural payload $\tau$ for non-duplicating ties.
Step vs SafeStep (artifact scope)
For the SafeStep relation, a single unified computable triple lex order $(\delta,\kappa^{M},\tau)$ suffices and is the canonical certificate used in this paper (Meta/ComputableMeasure.lean, with core relation definitions in Meta/SafeStep_Core.lean). For the unrestricted full relation Step, termination is not claimed in this artifact.
Theorem 4.1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}_{c}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}_{c}(t)$.
@@
DM and $\tau$ roles on rules (explicit)
For rec-zero, the $\kappa^{M}$ component provides a strict DM drop. For merge-cancel, the $\kappa^{M}$ component ties at 0, and the strict decrease is provided by the $\tau$ component. For rec-succ, the outer $\delta\text{-flag}$ drops from $1$ to 0. For eqW-rules in SafeStep (where $\kappa^{M}(a)=0$ is a guard), strict decrease is discharged in the computable $\tau$ component; the more general unguarded lemma in the artifact also includes a $\kappa^{M}(a)\neq 0$ branch discharged by DM.
  • DM premise: in duplicating-sensitive branches, the multiset of RHS pieces is strictly DM-smaller than the removed LHS redex profile.
  • $\tau$ premise: when the $\delta\text{-flag}$ and $\kappa^{M}$ tie, a direct natural-number inequality closes the branch.
@@
Lemma 4.2 (No fixed +k+k or boolean flag orients a generic duplicator).
@@
so there is no strict drop when $\rho(S)\geq 1$. The fix used here is the computable triple stack $(\delta\text{-flag},\kappa^{M},\tau)$: DM handles multiset-sensitive branches and $\tau$ resolves ties.
Corollary 4.4 (Strong normalization).
@@
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\mathrm{delta}$ structures from their contexts.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 4.2).
@@
The final artifact path is the computable lex stack $\mu^{3}_{c}=(\delta,\kappa^{M},\tau)$, where DM handles multiset-sensitive branches and $\tau$ resolves tie cases.
@@ 5 A certified normalizer
A normalization function is defined by well-founded recursion on SafeStepRev, with well-foundedness certified by the computable measure stack $\mu^{3}_{c}$. The normalizer operates on the root step relation SafeStep (not the context closure SafeStepCtx); a term is in normal form when no root safe rule applies ($\mathrm{NormalFormSafe}(t)\iff\neg\exists u.\;\mathrm{SafeStep}\;t\;u$). Properties proven in the formalization include:
(Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
(Soundness) $\forall t.~\mathrm{SafeStepStar}(t,\mathrm{normalizeSafe}(t))\land\mathrm{NormalFormSafe}(\mathrm{normalizeSafe}(t)).$
@@ 6 Local confluence and Newman (guarded safe relation)
The Newman engine is formalized and the global local-join hypothesis ($\mathsf{locAll}:\forall\,a,\,\mathrm{LocalJoinAt}\;a$) is fully discharged in the artifact. Combined with Cor. 4.4, Newman’s Lemma [11] yields:
Theorem 6.1 (Confluence and unique normal forms).
@@
Instantiation (SafeStep). The formalization provides a comprehensive library of local-join lemmas over safe root/context shapes, fully discharging the $\mathsf{locAll}$ hypothesis to prove unconditional confluence. In contrast, the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$; thus unrestricted full-step confluence is not claimed. The SafeStep relation introduces guard conditions that avoid this divergence pattern.
@@
  • SN (SafeStep). The computable KO7 triple measure $\mu^{3}_{c}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local-Join Library (SafeStep). Local-join lemmas per root shape and context wrappers are provided; the global hypothesis $\mathsf{locAll}$ is discharged in the artifact (locAll_safe), yielding unconditional confluence and unique NFs.
@@
The following table records representative safe root configurations with explicit local-join lemmas. We also exhibit a full-step non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ and discharge many guarded eqW cases via context wrappers. Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
@@
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ localJoin_all_safe (eqW/refl branch; target void)
For the $\mathrm{eqW}\,a\,a$ row, the relevant branch inside localJoin_all_safe is discharged via localJoin_of_unique.
@@
Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$. Here $\mathrm{deltaFlag}:\mathrm{Trace}\to\mathbb{N}$ is a binary root-shape detector for $\mathrm{rec}\Delta\,\_\,\_\,(\mathrm{delta}\,\_)$ defined in the artifact, so the predicate is decidable. Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
@@
$\mathrm{rec}\Delta\,b\,s\,(\mathrm{delta}\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\Delta\,b\,s\,n)$$\delta\text{-flag}$ drop
7 Impossibility results: The Fourteen Failed Methods
We document the failure of fourteen distinct termination strategy families, motivating the triple-lexicographic measure with Dershowitz-Manna orderings used in the safe fragment. Twelve of these failures are not merely heuristic; they are formally witnessed by machine-checked counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Conjecture_Boundary.lean). The remaining two (local helper lemmas and path orders) are treated as meta-theoretical and methodological barriers.